7k+6/3k+4 = 3k+4/2k+1 solve the equation for the k
Answer by Cromlix(4381)

Hi there,
7k+6/3k+4 = 3k+4/2k+1
Using a common denominator
(3k + 4)(2k + 1)
Cross multiply
(7k + 6)(2k + 1)/(3k + 4)(2k + 1)= (3k + 40(3k + 40/(3k + 4)(2k + 1)
Ignoring the denominator now you have cross multiplied.
Multiply out numerator brackets.
14k^2 + 19k + 6 = 9k^2 + 24k + 16
Collect like terms
14k^2 - 9k^2 + 19k - 24k + 6 - 16 = 0
5k^2 - 5k - 10 = 0
Divide throughout by 5
(k^2 - k - 2)= 0
Factorise
(k + 1)(k - 2) = 0
k + 1 = 0
k = -1
k - 2 = 0
k = 2
So, k = -1 or 2
Hope this helps:-)